OpenStudy (anonymous):

Calculate the area of triangle ABC with altitude CD, given A (-7, -1), B (-1, 5), C (0, 0), and D (-3, 3). (5 points) Select one: a. 9 square units b. 18 square units c. 18.5 square units d. 21 square units

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

yea I did

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

Sorry, wait a minute.

OpenStudy (anonymous):

Find the length of CD and the length of AB and the area will 1/2 CDx AB

OpenStudy (anonymous):

I'm guessing C.

OpenStudy (anonymous):

\[ CD= 3 \sqrt 2\\ AB = 6 \sqrt 2 Area = \frac 12 (3 \sqrt 2)(6 \sqrt 2)=18 \]

OpenStudy (anonymous):

so b then rigth

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[ CD= 3 \sqrt 2\\ AB = 6 \sqrt 2\\ Area = \frac 12 (3 \sqrt 2)(6 \sqrt 2)=18 \]

OpenStudy (anonymous):

Did you understand it?

OpenStudy (anonymous):

yes I di thank you

OpenStudy (anonymous):

YW